1) Using the Z table, what proportion of scores are Below a Z score of 1.50?

2) Using the Z table, what Z scores would you have if 30% of the data lies below you but above the mean?

3 Using the Z table, what proportion of scores are between the mean and a Z score of 1.50?

4) Using the Z table, what proportion of scores are above a Z score of 1.50?

(plz show steps, thanks)
Answer #1

1)

P(Z < 1.50) = 0.9332 (From Z table)

2)

We have to calculate z such that P(0 < Z < z) = 0.30

P(Z < z) - P(Z < 0) = 0.30

P(Z < z) - 0.5 = 0.30

P(Z < z) = 0.80

From Z table, z-score for the probability of 0.80 is 0.8416

z = 0.84

3)

P(0 < Z < 1.50) = P(Z < 1.50) - P(Z < 0)

= 0.9332 - 0.50

= 0.4332

4)

P(Z > 1.50) = 1 - P(Z < 1.50)

= 1 - 0.9332

= 0.0668